Dark current of the HQE PD and other PDs were measured.
- The HQE PDs were loaded on the new PD transportation cages (Attachment 1)
The PDs are always shorted by a clean PD plugs. The PD element is still capped with Kapton seals.
- The assignment of the container/slot and the PDs are as follows
- The measurement has been done with KEITHLEY sourcemeter SMU2450.
- The result is shown in Attachment 2. Most of the PDs show the dark current of ~3nA at 15V bias. C1-05 and C1-07 showed higher dark current at high V region. We should avoid using them for the aLIGO purpose. I hope they are still OK at low bias V if there is no noise issue (TBC). You can not read the PD names on the plot for the nominal ones, but that's OK as they are almost equivalent.
- As a comparison, the dark current of a C30655 (serial #10) was measured. Considering a DC current due to an anbient light (although the PD was covered), the dark current of the HQE PD seems double of C30655.
- Taking an advantage of having the setup, I took the same measurement for the Laser Comp. PDs in ATF. I gave the identification as #1 and #2. #1 has full-length legs while #2 has trancated legs. As Zach reported before, they showed significantly high dark current. (Attachment 3)